OpenStudy (anonymous):

What is the range of f(x) = ^-x? x≥ 0 y ≥ 0 y ≤ 0 all real numbers

OpenStudy (bobo-i-bo):

Do you know that the definition of the range of a function is?

OpenStudy (anonymous):

No..

OpenStudy (welshfella):

something is missing in that expression . What is taken to the power -x ?

OpenStudy (anonymous):

\[f(x)=\sqrt{-x}\]

OpenStudy (anonymous):

That's all there is.

OpenStudy (bobo-i-bo):

The range of a function is all the numbers which you can get as output from the function. For example, if the function was f(x)=3x, then since f(3)=3*3=9, 9 is a possible output you can get from the function, so 9 is part of the range of the function. So the question is, for your given function, what is the set of possible outputs which you can get from your function. Get what I mean? :)

OpenStudy (anonymous):

Would it be \[y \ge0\] then?

OpenStudy (bobo-i-bo):

yup!

OpenStudy (anonymous):

Thank you! c:

OpenStudy (bobo-i-bo):

np :)

OpenStudy (welshfella):

y ≤ 0 because if x > 0 then we would have the square root of a negative number for which is no real number solution.